OpenStudy (anonymous):

How to factor the expression completely: ((1+(4/x))^2)+((1-(4/x))^2)

OpenStudy (mertsj):

\[(1+(\frac{4}{x})^2)+(1-(\frac{4}{x})^2)\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

Actually the (^2) is on the outside of the parentheses.

OpenStudy (mertsj):

\[(1+\frac{4}{x})^2+(1-\frac{4}{x})^2\]

OpenStudy (mertsj):

Is that it?

OpenStudy (anonymous):

Yes! Sorry for the confusion..

OpenStudy (mertsj):

Expand each binomial and collect the like terms.

OpenStudy (anonymous):

Does that simplify to 16/x?

OpenStudy (anonymous):

I am so sorry again, but there's supposed to be a minus between the two quantities.. I got mixed up after all the parentheses

OpenStudy (mertsj):

\[1+\frac{8}{x}+\frac{16}{x^2}+1-\frac{8}{x}+\frac{16}{x^2}\]

OpenStudy (mertsj):

Collect the like terms.

OpenStudy (anonymous):

With the minus in between I get 16/x? Is that the answer then?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

Thank you so much!

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

Could you help with another problem?

OpenStudy (mertsj):

Only if you double check to make sure you have posted it correctly before I work on it.

OpenStudy (anonymous):

I will. I'll attach the file, thank you.

OpenStudy (anonymous):

The problem is in the attachment

OpenStudy (mertsj):

Factor out: \[(\frac{1}{2}x ^{-\frac{1}{2}})(7x+4)^{-\frac{1}{2}}\]

OpenStudy (anonymous):

That's the first portion of the problem right?

OpenStudy (mertsj):

It says to factor completely. That is what I would do as a first step is factor out those two common factors.

OpenStudy (anonymous):

How do I factor out the (7x+4)^(-1/2) if the first one is positive?

OpenStudy (mertsj):